OpenStudy (anonymous):

prove the identity: cos+tan^2cos=1/cos

OpenStudy (isaiah.feynman):

Do you have the original question?

OpenStudy (anonymous):

that is the original question. I have to prove that one side is equal to the other.

OpenStudy (isaiah.feynman):

Start by reducing tan to sin and cos.

OpenStudy (unklerhaukus):

LHS: cos+tan^2cos = cos(1+tan^2) = cos(sec^2) = cos(1/cos^2) = 1/cos : RHS

OpenStudy (unklerhaukus):

from cos^2+sin^2 = 1 1+sin^2/cos^2 = 1/cos^2 1+tan^2 = sec^2